The trial balance before adjustment for Sheridan Company shows the following balances.

Dr. Cr.
Accounts receivable $74,000
Allowance for doubtful accounts 1,800
Sales revenue $430,000


The following cases are independent:

1. To obtain cash, Sheridan factors without recourse $22,000 of receivables with Anila Finance. The finance charge is 10% of the amount factored.
2. To obtain a one-year loan of $67,000, Sheridan assigns $77,000 of specific accounts receivable to Ruddin Financial. The finance charge is 6% of the loan; the cash is received.
3. The company wants to maintain the Allowance for Doubtful Accounts at 4% of gross accounts receivable.
4. The company wishes to increase the allowance account by 1.4% of sales.


Using the data above, prepare the journal entries to record each of the above cases. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

No.

Account Titles and Explanation

Debit

Credit

1.

2.

3.

4.
Answer #1
No Account Titles and Explanation Debit Credit
1 Cash 19800
Loss on sale of Receivables 2200 =22000*10%
     Accounts receivable 22000
2 Cash 62980
Interest expense 4020 =67000*6%
      Notes payable 67000
3 Bad Debts expense 4760 =(74000*4%)+1800
     Allowance for doubtful accounts 4760
4 Bad Debts expense 6020 =430000*1.4%
     Allowance for doubtful accounts 6020